# BUFFER-X: Towards Zero-Shot Point Cloud Registration in Diverse Scenes

### v1

@@ E Qualitative results in diverse scenes
Fig. A5: Qualitative results on indoor point cloud registration (T-B): ScanNet++i, ScanNet++F, and TIERS sequences. (a) Input source (yellow) and target (cyan) point clouds before registration. (b) Registration results obtained using our BUFFER-X. (c) Ground truth alignment. Visualization demonstrates that BUFFER-X achieves accurate alignment, closely matching the ground truth.
@@
Fig. A6: Qualitative results on outdoor point cloud registration: (T-B): Oxford, KAIST, and MIT sequences. (a) Input source (yellow) and target (blue) point clouds before registration. (b) Registration results obtained using our BUFFER-X. (c) Ground truth alignment. Visualization demonstrates that BUFFER-X achieves accurate alignment, closely matching the ground truth.
Therefore, these qualitative results further support the zero-shot registration capability of our BUFFER-X, regardless of the environment, sensor type, acquisition setup, or range.
Env. Indoor Outdoor Average rank
Conven- tional FPFH [63] + FGR [94] + [Uncaptioned image] 62.53 15.42 77.68 92.31 80.60 98.74 100.00 89.80 74.78 91.87 99.00 8.55
FPFH [63] + Quatro [43] + [Uncaptioned image] 8.22 1.74 9.88 97.27 86.57 99.10 100.00 91.46 79.57 51.05 91.03 10.82
FPFH [63] + TEASER++ [77] + [Uncaptioned image] 52.00 13.25 66.15 97.22 73.13 98.92 100.00 89.20 71.30 93.69 99.34 8.82
Deep FCGF [22] 8.04 0.17 19.96 23.07 77.82 98.92 95.38 88.34 82.17 6.59 75.08 14.82
Predator [32] N/A (Err) N/A (Err) N/A (Err) N/A (Err) 69.43 99.82 100.00 N/A (OOM) 54.78 55.68 89.04 14.09
GeoTransformer [59] N/A (Err) N/A (Err) N/A (Err) N/A (Err) N/A (Err) 99.82 100.00 63.84 93.91 77.00 73.42 12.73
BUFFER [5] N/A (Err) N/A (Err) 17.60 88.84 93.34 99.64 100.00 99.50 95.22 98.18 99.34 8.27
PARENet [80] N/A (Err) N/A (Err) N/A (Err) N/A (Err) N/A (Err) 99.82 97.69 57.51 75.22 68.30 66.11 15.82
Table A3: Quantitative performance comparison on rate to evaluate generalization capability. All the models were trained only on KITTI [28] and RANSAC was used with the maximum iteration of 50K. The icons represent oracle tuning ([Uncaptioned image]) to set proper voxel size and search radius to each dataset, and scale-up alignment ([Uncaptioned image]) to normalize dataset scales to be similar to that of the KITTI data (e.g., the scale of 3DMatch, which typically uses a voxel size of 0.025, is adjusted to match the scale of KITTI, which typically uses a voxel size of 0.3, by multiplying by $\frac{0.3}{0.025}$). N/A (Err) indicates that the network cannot run due to an insufficient number of points remaining after voxelization with the voxel size typically used for outdoor settings, making keypoint extraction or descriptor generation infeasible. N/A (OOM) indicates an out-of-memory error caused by excessive memory usage.
F Quantitative results using KITTI
Therefore, while training on KITTI led to some variations in overall performance, our approach remains more robust than other state-of-the-art methods in zero-shot registration.

### v3

@@ F Qualitative results in diverse scenes
Fig. A5: Qualitative results on indoor point cloud registration (T-B): ScanNet++i, ScanNet++F, and TIERS sequences. (a) Input source (yellow) and target (cyan) point clouds before registration. (b) Registration results obtained using our BUFFER-X, trained only on 3DMatch. (c) Ground truth alignment. Visualization demonstrates that BUFFER-X achieves accurate alignment, closely matching the ground truth.
@@
Fig. A6: Qualitative results on outdoor point cloud registration: (T-B): Oxford, KAIST, and MIT sequences. (a) Input source (yellow) and target (blue) point clouds before registration. (b) Registration results obtained using our BUFFER-X, trained only on 3DMatch. (c) Ground truth alignment. Visualization demonstrates that BUFFER-X achieves accurate alignment, closely matching the ground truth.
Therefore, these results further support the zero-shot registration capability of our BUFFER-X, regardless of the environment, sensor type, acquisition setup, or range.